# DUDF: Differentiable Unsigned Distance Fields with Hyperbolic Scaling

### v1

@@ 4.5 Ablation study
We now dissect the impact of individual loss components on the network’s accuracy through an ablation study. The methodology involves selectively deactivating terms in our loss function (Eq. 4) to isolate their contributions. We adopted the experimental framework from the prior section, training on 30 meshes from the DeepFashion dataset [48]. To elucidate the role of boundary conditions in enhancing reconstruction quality, we incorporated a normal consistency metric into our analysis. This metric is computed as the mean value of the absolute cosine similarities between each point to its closest neighbor in the other point cloud. We report metrics for reconstructions performed using MC2 [14], yielding results analogous to those obtained with MC1 [46]. Additionally, we contrasted the effects of approximating the distance function $d_{\mathcal{S}}$ with $t_{\mathcal{S}}$ (Eq. 1), utilizing sine and $ReLU$ activation functions (each with distinct Eikonal problems).
The findings presented in Table 3 underscore the significance of both the Eikonal equation (Eq. 5) and the $MCurv$ boundary condition (Eq. 8) in achieving superior normal consistency, which notably influences the rendering quality. Moreover, the refinement step during training is an important element for normal consistency improvement, offering substantial increased performance for a small additional computational cost.
@@
Table 3: Quantitative impact of each loss component on the reconstruction accuracy of our network. We report training time, L1, and L2 mean Chamfer distances ×$\times 10^{3}$, and Normal Consistency (NC).
@@ 5 Limitations and future work
Our method encounters certain limitations. Primarily, by approximating the function $t_{\mathcal{S}}$ (Eq. 1), the Eikonal equation requires supervision with ground truth unsigned distances, in contrast to signed distance approaches where it is constant everywhere. In addition, the quadratic behavior of the function near the isosurface produces a wider near-zero strip which can lead to slightly inflated representations. This effect can be lessened by computing square roots, at the expense of an increased computational cost. Secondly, we observed that surfaces extracted using gradient-based Marching Cubes methods can sometimes be less smooth than directly rendering the function using sphere tracing. This is likely due to the current reconstruction algorithm’s sensitivity to noise, an issue that might be addressed by employing a tangency-aware surface reconstruction strategy [33]. Additionally, computing the maximum curvature direction field is computationally expensive, requiring three network passes and the diagonalization of the Hessian matrix. However, future improvements in the network’s design could reduce the number of parameters, thus lessening the computational burden associated with computing topological features.
@@ 6 Conclusion
In this work we introduced Differentiable Unsigned Distance Fields (DUDF) with Hyperbolic Scaling, a novel approach that addresses inherent limitations of traditional UDFs in representing open surfaces. By applying a hyperbolic transformation to the distance field, we define a new variant of the Eikonal problem, tailored with unique boundary conditions. This adaptation enables our model to capitalize on the robust framework of continuously differentiable neural networks, thus enhancing reconstruction quality and training efficiency. The conducted experiments provide evidence that our approach can lead to improvements in reconstruction quality and computational performance when compared to several state-of-the-art methods. Moreover, the ability to accurately calculate topological features such as normals and curvatures is a notable benefit of our model, offering additional utility in geometric processing tasks and rendering. While our results are promising, showing better accuracy and efficiency in most cases, we acknowledge the need for further research and development. The potential applications of DUDF in 3D geometry processing are broad, and we are optimistic about its utility in practical scenarios. Future work will focus on exploring these applications and enhancing the model’s performance.
@@ 7 Aknowledgments
This project was supported by grants from Salesforce, USA (Einstein AI 2020), National Scientific and Technical Research Council (CONICET), Argentina (PIP 2021-2023 GI -11220200102981CO), and Universidad Torcuato Di Tella, Argentina. We also thank Facundo Fainstein and Tomas Delgado for valuable discussions on the Eikonal problem formulation.

### v2

@@ 4.5 Ablation study
We now dissect the impact of individual loss components on the network’s accuracy through an ablation study. The methodology involves selectively deactivating terms in our loss function (Eq. 4) to isolate their contributions. We adopted the experimental framework from the prior section, training on 30 meshes from the DeepFashion dataset [48]. We report metrics for reconstructions performed using MC2 [14], yielding results analogous to those obtained with MC1 [46].
The findings presented in Table 3 underscore the significance of the Eikonal equation (Eq. 5) in achieving superior normal consistency ($\lambda_{g}$). Moreover, the refinement step during training ($\lambda_{\mu},\lambda_{\sigma}$) is an important element for normal consistency improvement, offering substantial increased performance for a small additional computational cost. Although $\mathcal{L}_{MCurv}$ (Eq. 8) does not offer substantial quantitative benefits ($\lambda_{g}$), we found that this loss term becomes very relevant qualitatively when rendering through direct methods (see Supp. Material for more details). Additionally, we contrasted the effects of approximating the ground truth distance function $d_{\mathcal{S}}$ (with its distinct Eikonal problem), using the same network architecture and sampling scheme, with sine and ReLU activations. The findings from this study show that such an approach leads to suboptimal outcomes, thereby underscoring the effectiveness of $t_{\mathcal{S}}$ and the derived Eikonal problem proposed.
Finally, we ablated parameter $\alpha$ (see Supp. Material). As $\alpha$ gets larger, function $t_{\mathcal{S}}$ closely approximates $d_{\mathcal{S}}$, increasing the reconstruction error (probably due to the non-differentiability at the isosurface). Smaller $\alpha$ values enlarge the quadratic strip near the isosurface, which is harder to supervise effectively and hinders the performance of MC.
@@
Table 3: Quantitative impact of each loss component on the reconstruction accuracy of our network. We report training time, L1, and L2 mean Chamfer distances ×$\times 10^{3}$, and Normal Consistency (NC). Last two rows correspond to approximating the ground truth distance function $d_{\mathcal{S}}$ instead of $t_{\mathcal{S}}$.
@@ 5 Limitations and future work
Our method encounters certain limitations. Primarily, by approximating the function $t_{\mathcal{S}}$ (Eq. 1), the Eikonal equation requires supervision with ground truth unsigned distances, in contrast to signed distance approaches where it is constant everywhere. In addition, the quadratic behavior of the function near the isosurface produces a wider near-zero strip which can lead to slightly inflated representations. This effect can be lessened by computing square roots, at the expense of an increased computational cost. Besides hyperbolic scaling, we experimented with the square distance and SmoothL1. The former requires direct rendering to compute square roots at every marching step and SmoothL1 does not yield the true distance for points far from the isosurface when $\delta$ is nonzero. This is not the case for our hyperbolic function, that satisfies $t_{\mathcal{S}}=d_{\mathcal{S}}$ for distant domain points. Future research can find other suited functions for distance field learning. Secondly, we observed that surfaces extracted using gradient-based Marching Cubes methods can sometimes be less smooth than directly rendering the function using sphere tracing. This is likely due to the current reconstruction algorithm’s sensitivity to noise, an issue that might be addressed by employing a tangency-aware surface reconstruction strategy [33]. Additionally, computing the maximum curvature direction field is computationally expensive, requiring three network passes and the diagonalization of the Hessian matrix. Future network’s designs could reduce the number of parameters, thus lessening the computational burden associated with computing topological features.
@@ 6 Conclusion
In this work we introduced Differentiable Unsigned Distance Fields (DUDF) with Hyperbolic Scaling, a novel approach that addresses inherent limitations of traditional UDFs in representing open surfaces. By applying a hyperbolic transformation to the distance field, we define a new variant of the Eikonal problem, tailored with unique boundary conditions. This adaptation enables our model to capitalize on the robust framework of continuously differentiable neural networks, thus enhancing reconstruction quality and training efficiency. The conducted experiments provide evidence that our approach can lead to improvements in reconstruction quality and computational performance when compared to several state-of-the-art methods. Moreover, the ability to accurately calculate topological features such as normals and curvatures is a notable benefit of our model, offering additional utility in geometric processing tasks and rendering. While our results are promising, showing better accuracy and efficiency in most cases, we acknowledge the need for further research and development. The potential applications of DUDF in 3D geometry processing are broad, and we are optimistic about its utility in practical scenarios.
@@ 7 Aknowledgments
This project was supported by grants from Salesforce, USA (EinsteinAI 2020), National Scientific and Technical Research Council (CONICET), (PIP 2021-2023 GI -11220200102981CO), and Universidad Torcuato Di Tella, Argentina. We also thank Facundo Fainstein and Tomas Delgado for valuable discussions on the Eikonal problem.